# Motivic Steenrod operations at the characteristic via infinite ramification

### v1

@@ 2. $(T_{\mathrm{cts}}^{n}(X_{\overline{k}};\mathbb{Z}_{p}))$
In the notation of Definition 4.22, one can formulate the crystalline Tate conjecture as
  1. ($T^{n}(X)$)
    The cycle class map $\mathrm{CH}^{n}(X)\otimes\mathbb{Q}\rightarrow H^{2n}_{\mathrm{crys}}(X/W(k))[\tfrac{1}{p}])^{\phi=p^{n}}$ is surjective.
By Remark 4.21, $T^{n}(X;\mathbb{Z}_{p})\Rightarrow T^{n}(X)$. On the other hand, a theorem of Morrow [Mor19, Proposition 4.1, Theorem 4.3] shows that if $T^{1}(X)$ is equivalent to the usual Tate conjecture in $\ell$-adic cohomology and furthermore that the veracity of $T^{1}(X)$ for surfaces implies it for all other smooth projective, geometrically connected varieties.
Remark 4.24.
We note that if $T^{n}(X_{k^{\prime}};\mathbb{Z}_{p})$ is true for all $k^{\prime}$, then $T_{\mathrm{cts}}^{n}(X_{\overline{k}};\mathbb{Z}_{p})$ is true. Hence, the most optimistic hope that one can ask for is the veracity of $T_{\mathrm{cts}}^{n}(X_{\overline{k}};\mathbb{Z}_{p})$. This story is similar to the $\ell$-adic, integral counterpart of the statement as discussed by Schoen in [Sch98]. We also make a remark that the formation of syntomic cohomology does not preserve filtered colimits of schemes even along affine transition maps. Therefore, there it is not true that the map $\mathrm{colim}_{k^{\prime}}\mathrm{CrysTate}_{n}(X_{k^{\prime}})\rightarrow\mathrm{CrysTate}_{n}(X_{\overline{k}})$ is an isomorphism. Nonetheless by Remark 4.21, the map $\mathrm{colim}_{k^{\prime}}\mathrm{CrysTate}_{n}(X_{k^{\prime}})\rightarrow(H^{2i}_{\mathrm{crys}}(X_{\overline{k}}/W(k))[\tfrac{1}{p}])^{\phi=p^{i}}$ is rationally surjective.
Example 4.25 ($T^{1}(X;\mathbb{Z}_{p})$ for smooth projective curves).
The first interesting case of the integral Tate conjecture is for smooth projective, geometrically connected curves over a finite field, $(T^{1}(X;\mathbb{Z}_{p}))$. In this case the conjecture asks that the map
$\mathrm{Pic}(X)\otimes\mathbb{Z}_{p}\rightarrow\mathrm{CrysTate}_{1}(X)$
is surjective. We claim that
@@
which verifies $T^{1}(X;\mathbb{Z}_{p})$. Indeed (for all qcqs scheme $X$) weight one syntomic cohomology is calculated as
@@
Example 4.26 ($T^{\dim(X)}(X;\mathbb{Z}_{p})$).
Let $X$ be a smooth projective, geometrically connected variety over a finite field of dimension $d$; we sketch the veracity of $T^{d}(X;\mathbb{Z}_{p})$. This result can be deduced using Kato–Saito’s unramified class field theory for smooth projective varieties over a finite field [KS83]; we also note that Gros [Gro85, Théorème 3.2.0] has a direct proof of this result using cohomological methods which are closed to this paper. This result also generalizes Example 4.25. To begin with, we have Milne’s duality [Mil86]:
@@
[…]
@@
Having verified some general cases of $T^{n}(X;\mathbb{Z}_{p})$, we now turn to counterexamples.
@@
We now provide a counterexample to the integral crystalline Tate conjecture in codimension two. The counterexample is a little different from the ones produced in [AH62, CTS10] and it involves the product of an approximation to classifying space with an ordinary elliptic curve. To this end, we review the calculation of mod-$p$ syntomic cohomology of elliptic curves over $\overline{\mathbb{F}}_{p}$.
@@
(24) 00{0}${H^{0}_{\mathrm{syn}}(E_{n};\mathbb{F}_{p}(0))}$${H^{0}(E_{n};\mathcal{O}_{E_{n}})}$${H^{0}(E_{n};\mathcal{O}_{E_{n}})}$${H^{1}_{\mathrm{syn}}(E_{n};\mathbb{F}_{p}(0))}$${H^{1}(E_{n};\mathcal{O}_{E_{n}})}$${H^{1}(E_{n};\mathcal{O}_{E_{n}})}$${H^{2}_{\mathrm{syn}}(E_{n};\mathbb{F}_{p}(0))}$${0,}$$\scriptstyle{1-F}$$\scriptstyle{\delta_{1}}$$\scriptstyle{1-F}$$\scriptstyle{\delta_{2}}$
@@
(25) 00{0}${H^{1}_{\mathrm{syn}}(E_{n};\mathbb{F}_{p}(1))}$${H^{0}(E_{n};\Omega^{1}_{E_{n}})}$${H^{0}(E_{n};\Omega^{1}_{E_{n}})}$${H^{2}_{\mathrm{syn}}(E_{n};\mathbb{F}_{p}(1))}$${H^{1}(E_{n};\Omega^{1}_{E_{n}})}$${H^{1}(E_{n};\Omega^{1}_{E_{n}})}$${H^{3}_{\mathrm{syn}}(E_{n};\mathbb{F}_{p}(1))}$${0.}$$\scriptstyle{1-C}$$\scriptstyle{\delta_{3}}$$\scriptstyle{1-C}$$\scriptstyle{\delta_{4}}$
@@
    If $E$ is supersingular, i.e., the action of $F$ on $H^{1}(E;\mathcal{O}_{E})$ is zero, then $1-F$ acts as the identity on $H^{1}(E_{n};\mathcal{O}_{E_{n}})$. Combining this observation with Milne’s duality result, we observe that the syntomic cohomology of $E_{n}$ is generated by the classes $1,\epsilon_{n},y,$ and $\tau_{n}$.
@@
    By Milne’s duality result, we observe that the operator $1-C$ on $H^{0}(E_{n};\Omega^{1}_{E_{n}})$ has a one-dimensional kernel and cokernel. Let $z\in H^{1}(E_{n};\mathbb{F}_{p}(1))$ be the generator of the kernel, and let $\rho_{n}$ be the generator of the image of $\delta_{3}$.
@@
To finish our analysis, we observe that the connecting maps $\delta_{i}$ vanish in the colimit because the operators $1-F$ and $1-C$ induce surjective maps on cohomology in the colimit. This can be seen by a computation using the Frobenius linearity of $F$ and $C$ on the respective coherent cohomology groups, and the fact that these groups are one-dimensional over $\overline{\mathbb{F}}_{p}$. Thus, the generators of the syntomic cohomology of $E$, i.e., the classes that survive the colimit, are
@@
There exists a smooth, projective $\mathbb{F}_{p}$-variety $X$ such that $T_{\mathrm{cts}}^{2}(X_{\overline{\mathbb{F}_{p}}};\mathbb{Z}_{p})$ does not hold. In other words, there is a class in $\mathrm{colim}_{\mathbb{F}_{p}\subset k^{\prime}\subset\overline{k}}\mathrm{CrysTate}_{2}(X_{k^{\prime}})$ which does not come from an algebraic cycle.
@@
At the prime $p=2$, the counterexample of Atiyah–Hirzebruch [AH62] is of dimension $7$. Ours is slightly of lower dimension since it is of dimension $2(2)+1=5$. More generally, for any prime $p$ if it is of dimension $12p+$ 1. We expect something like [SV05, Theorem 1] to also hold in characteristic $p>0$ where the minimal dimension counterexample detected via Steenrod operations is at least $p$. It would be interesting to find counterexamples of lower dimension.
@@ 4.4. The Wu formula for motivic and syntomic cohomology
where $\mathcal{E}$ is a locally free rank $r$ sheaf on $X\in\mathrm{Sm}^{?}_{S}$. If $E\in\mathrm{CAlg}(\mathrm{h}\mathrm{MS}^{?}_{S})$, then $t(\mathcal{E})$ is furthermore implemented by multiplication with the Thom class $t(\mathcal{E})\in E^{2r,r}(\mathrm{Th}_{X}(\mathcal{E}))$, which is the image of $1\in E^{0,0}(X)$ under the isomorphism (28).
@@
in $\mathrm{MS}^{?}_{S}$. If $E\in\mathrm{CAlg}(\mathrm{h}\mathrm{MS}^{?}_{S})$ is oriented, then we can compose this with the Thom isomorphism which takes the form
@@
where $i$ is a quasi-smooth closed immersion and $p$ is the projection map. The first author and Shin has defined, for $E\in\mathrm{CAlg}(\mathrm{h}\mathrm{MS}^{?}_{S})$ a pushforward map
@@
Let $E\in\mathrm{CAlg}(\mathrm{h}\mathrm{MS}^{?}_{S})$ be oriented. Then, the Gysin pushforwards satisfy the following properties:
@@
    Naturality: If $\varphi:E\to F$ is an orientation-preserving map of oriented theories in $\mathrm{CAlg}(\mathrm{h}\mathrm{MS}^{?}_{S})$, then the squares
@@
be a map in $\mathrm{CAlg}(\mathrm{h}\mathrm{MS}^{?}_{S})$, and let $e$ and $f$ be orientations of $E$, and $F$ respectively. Evaluating $\varphi$ on $\mathbb{P}^{\infty}_{S,+}$ induces a map of bigraded rings
@@
where $A=G(c,x)$ for some formal power series $G(a,b)\in E^{\star,\star}[[a,b]]$. As $c-x$ annihilates $x$, we see that $c_{1}(\mathscr{L}(-1))=c-x$, which is exactly the desired formula for a line bundle. ∎
@@
Let $\varphi\colon E\to F$ be a possibly non-orientation-preserving map in $\mathrm{CAlg}(\rm{h}\mathrm{MS}^{?}_{S})$. Let $f\colon X\to Y$ be a projective [^18] Here, a projective morphism is a map that factors through a closed embedding $i\colon X\hookrightarrow\mathbb{P}^{n}_{Y}$ for some $n\geq 0$. quasi-smooth morphism in $\mathrm{Sm}^{?}_{S}$, Then, for all $x\in E^{\star,\star}(X)$,
@@ Appendix A Approximation of syntomic cohomology of classifying stacks
As in the proof of [ABM21, Theorem 1.2], we can find a $k$-linear representation $V$, a $d$-dimensional complete intersection $Z\subset\mathbb{P}(V)$ such that $Z$ is $G$-stable, $G$ acts freely on $Z$ and $Z/G\simeq[Z/G]$ is smooth and projective. For such a $Z$, [ABM21, Proposition 5.3 and 5.10] implies that the induced map $[Z/G]\simeq Z/G\rightarrow[\mathbb{P}(V)/G]$ is a Hodge $d$-equivalence. Therefore, by Lemma A.2, the induced map on syntomic cohomology has cofibres in degrees $\geq d$. The claim then follows from the projective bundle formula in syntomic cohomology; we remark that this result for schemes is [BL22, Theorem 9.1.1] and the result follows by Kan extension for stacks which, in particular, shows that $H^{i}_{\mathrm{syn}}(\mathrm{B}G;\mathbb{F}_{p}(j))$ is a summand of $H^{i}_{\mathrm{syn}}([\mathbb{P}(V)/G];\mathbb{F}_{p}(j))$. Now, noting that $\mathrm{Spec}(k)\rightarrow\mathrm{B}G$ is an fpqc surjection, which pulls-back along $[\mathbb{P}(V)/G]\rightarrow\mathrm{B}G$ to the fpqc surjection $\mathbb{P}(V)\rightarrow[\mathbb{P}(V)/G]$ the projective bundle formula follows from the one for schemes by the formula (35). ∎
@@ Appendix B Calculations in syntomic and motivic cohomology of certain classifying stacks
Throughout this section, we fix an perfect field $k$ of characteristic $p>0$. For an algebraic stack $\mathfrak{X}$ we define $\mathbb{F}_{p}(j)^{?}(\mathfrak{X})$ where ?$=\mathrm{mot}$ or $\mathrm{syn}$ via right Kan extension from schemes. We write
@@
the bigraded cohomology ring of $k$ where ?$=\mathrm{syn}$ or $\mathrm{mot}$. As already recorded in Example 4.11 we know how the cohomology of $B\mu_{p}^{\times s}$ takes the form:
@@
Here, |||-|| - | refers to the topological/simplicial degree of a class/operation.
@@
We first give references for these equalities in the characteristic zero field $\mathbb{K}$. If $p$ is odd, then both equalities is [HW19, Lemma 13.11] (note that there should actually be a sign in the formulas because it is the case in topology and op. cit. obtained these formulas from the topological Steenrod algebra); note that in [HW19, Definition 13.10] the $\mathrm{Q}_{n}$’s are defined inductively by the right hand side of (1) and the cited reference proves the equality with our definition of $\mathrm{Q}_{n}$. If $p=2$, then we note that [Voe03b, Example 13.7] cautions that the commutator formula might not hold; it might also not be a derivation by [HW19, Remark 13.4.1]. However, the failure of both equalities are due to the presence of the element $\rho$ in motivic cohomology; such an element is zero if $-1$ is a sum of squares in the base field. We note that by examining [HW19, Corollary 13.14, Remark 13.14.1] both equalities hold in the case of $p=2$ modulo $\rho$. By the construction of our operations in characteristic $p>0$, both equalities now hold for all primes $p$ since $\rho$ is zero in characteristic $p>0$. The result for syntomic cohomology follows immediately. ∎

### v2

@@ 3. $(C^{n}_{X_{k}})$
First we note that, after tensoring with $\mathbb{Q}_{p}$, the statements $(A^{n}_{X}),(B^{n}_{X})$ and $(C^{n}_{X})$ are equivalent by a standard transfer argument. These are in turn, thanks to the isomorphisms of Eq. (24), equivalent to the crystalline Tate conjecture, which is the statement that:
  1. ($\dagger^{n}_{X}$)
    The cycle class map $\mathrm{CH}^{n}(X)\otimes\mathbb{Q}_{p}\rightarrow H^{2n}_{\mathrm{crys}}(X/W(k))[\tfrac{1}{p}])^{\phi=p^{n}}$ is surjective.
The reader should also note the following relationship between the crystalline Tate conjecture and the usual $\ell$-adic Tate conjecture: a theorem of Morrow (also proved by de Jong) [Mor19, Proposition 4.1, Theorem 4.3] shows that $\dagger^{1}_{X}$ is equivalent to the usual Tate conjecture in $\ell$-adic cohomology for divisors, and furthermore that the veracity of $\dagger^{1}_{X}$ for surfaces implies it for all other smooth projective, geometrically connected varieties.
Remark 4.27 ($(A^{n}_{X})\Rightarrow(B^{n}_{X})$).
Thanks to the surjection (23), evidently $(A^{n}_{X})\Rightarrow(B_{n}^{X})$. However, we do not expect the converse to hold as in the case of the $\ell$-adic integral Tate conjecture thanks to the examples of Scavia–Suzuki [SS24, Theorem 1.4].
Remark 4.28 ($(B^{n}_{X_{k^{\prime}}})\Rightarrow(C^{n}_{X}),k^{\prime}\gg k$).
We note that if $B^{n}_{X_{k^{\prime}}}$ is true for all finite extensions $k^{\prime}$ of $k$ , then $C^{n}_{X}$ is true. In fact, we only need the veracity of $B^{n}_{X_{k^{\prime}}}$ for all $k^{\prime}$ with enough elements. The statement $C^{n}_{X}$ is the most optimistic one. This story is similar to the $\ell$-adic, integral counterpart of the statement as discussed by Schoen in [Sch98]; the target in $C^{n}_{X}$ is the $p$-adic analog of the group:
$\bigcup_{U\leq G,\text{open}}H^{2n}(X_{\overline{k}},\mathbb{Z}_{\ell}(n))^{U},$
which is a target of the cycle class map studied in op. cit.. We also make a remark that the formation of syntomic cohomology does not preserve filtered colimits of schemes even along affine transition maps. Therefore, there it is not true that the map $\mathrm{colim}_{k^{\prime}}\mathrm{Tate}_{p}^{n}(X_{k^{\prime}})\rightarrow\mathrm{Tate}_{p}^{n}(X_{\overline{k}})$ is an isomorphism. Nonetheless by Remark 4.24, the map $\mathrm{colim}_{k^{\prime}}\mathrm{Tate}_{p}^{n}(X_{k^{\prime}})\rightarrow(H^{2i}_{\mathrm{crys}}(X_{\overline{k}}/W(k))[\tfrac{1}{p}])^{\phi=p^{i}}$ is rationally surjective.
Remark 4.29 (Variant with the syntomic cycle class map).
We can formulate the following version of $A^{n}_{X}$:
  1. ($A^{n}_{X,\mathrm{syn}}$)
    The syntomic cycle class map $\mathrm{CH}^{n}(X)\xrightarrow{\mathrm{cyc}^{n}_{\mathrm{syn}}}H^{2n}_{\mathrm{syn}}(X,\mathbb{Z}_{p}(n))$ is surjective.
Evidently, $A^{n}_{X,\mathrm{syn}}\Rightarrow A^{n}_{X}$. However, $A^{n}_{X}$ can still hold without $A^{n}_{X,\mathrm{syn}}$; non-algebraizable classes in syntomic cohomology could potentially vanish in crystalline cohomology. While we do not have an example of this phenomenon, we still propose $A^{n}_{X}$ as a better variant of the integral Tate conjecture which has a better chance of being true (and therefore harder to give counterexamples to) than $A^{n}_{X,\mathrm{syn}}$, though we anticipate that whenever $A^{n}_{X}$ holds it is only because one can prove $A^{n}_{X,\mathrm{syn}}$. Similar, “syntomic versions” of $B^{n}_{X}$ and $C^{n}_{X}$ can also be formulated and left to the reader.
Example 4.30 ($A^{1}_{X}$ for smooth projective curves).
The first interesting case of the integral Tate conjecture is for smooth projective, geometrically connected curves over a finite field, $(A^{1}_{X})$. In this case the conjecture asks that the map
$\mathrm{Pic}(X)\otimes\mathbb{Z}_{p}\rightarrow\mathrm{Tate}_{p}^{1}(X)$
is surjective. We claim a stronger statement, namely that
@@
which verifies $A^{1}_{X}$. Indeed (for all qcqs scheme $X$) weight one syntomic cohomology is calculated as
@@
Example 4.31 ($A_{X}^{\dim(X)}$).
Let $X$ be a smooth projective, geometrically connected variety over a finite field of dimension $d$; we sketch the veracity of $A^{d}_{X}$. This result can be deduced using Kato–Saito’s unramified class field theory for smooth projective varieties over a finite field [KS83]; we also note that Gros [Gro85, Théorème 3.2.0] has a direct proof of this result using cohomological methods which are closed to this paper. This result also generalizes Example 4.30. To begin with, we have Milne’s duality [Mil86]:
@@
$\mathrm{CH}_{0}(X)\otimes\mathbb{Z}_{p}$$H^{2d}_{\mathrm{syn}}(X;\mathbb{Z}_{p}(d))$$\lim\mathrm{CH}_{0}(X)/p^{r}$$\lim\pi_{1}^{\mathrm{ab}}(X)/p^{r}$$\scriptstyle{\mathrm{cyc}^{d}_{\mathrm{syn}}}$
@@
Having verified some general cases of $A^{n}_{X}$, we now turn to counterexamples. The mechanism by which this happens is via Steenrod operations as made precise by the following.
@@
The first systematic study of this conjecture and its counterexamples was first produced carefully in [CTS10]. Other counterexamples to the $\ell$-adic integral Tate conjecture have been studied by various authors including (but not limited to): [PY15, Ant16, Qui11, Ben25]. Our counterexample is in codimension two but is a little different from the ones produced in [AH62, CTS10] and it involves the product of an approximation to classifying space with an ordinary elliptic curve. To this end, we review the calculation of mod-$p$ syntomic cohomology of elliptic curves over $\overline{\mathbb{F}}_{p}$.
@@
[…]
@@
[…]
@@
    If $E$ is supersingular, i.e., the action of $F$ on $H^{1}(E;\mathcal{O}_{E})$ is zero, then 1F1-F acts as the identity on $H^{1}(E_{n};\mathcal{O}_{E_{n}})$. Combining this observation with Milne’s duality result, we observe that the syntomic cohomology of $E_{n}$ is generated by the classes $1,\epsilon_{n},y,$ and $\tau_{n}$.
@@
    By Milne’s duality result, we observe that the operator 1C1-C on $H^{0}(E_{n};\Omega^{1}_{E_{n}})$ has a one-dimensional kernel and cokernel. Let $z\in H^{1}(E_{n};\mathbb{F}_{p}(1))$ be the generator of the kernel, and let $\rho_{n}$ be the generator of the image of $\delta_{3}$.
@@
To finish our analysis, we observe that the connecting maps $\delta_{i}$ vanish in the colimit because the operators 1F1-F and 1C1-C induce surjective maps on cohomology in the colimit. This can be seen by a computation using the Frobenius linearity of $F$ and $C$ on the respective coherent cohomology groups, and the fact that these groups are one-dimensional over $\overline{\mathbb{F}}_{p}$. Thus, the generators of the syntomic cohomology of $E$, i.e., the classes that survive the colimit, are
@@
There exists a smooth, projective $\mathbb{F}_{p}$-variety $X$ such that $(C^{2}_{X})$ does not hold.
@@
At the prime $p=2$, the counterexample of Atiyah–Hirzebruch [AH62] is of dimension $7$. Ours is slightly of lower dimension since it is of dimension $2(2)+1=5$. More generally, for any prime $p$ if it is of dimension $2p+1$. We expect something like [SV05, Theorem 1] to also hold in characteristic $p>0$ where the minimal dimension counterexample detected via Steenrod operations is at least $p$. It would be interesting to find counterexamples of lower dimension.
@@ 4.4. The Wu formula for motivic and syntomic cohomology
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
where $A=G(c,x)$ for some formal power series $G(a,b)\in E^{\star,\star}[[a,b]]$. As cxc-x annihilates $x$, we see that $c_{1}(\mathscr{L}(-1))=c-x$, which is exactly the desired formula for a line bundle. ∎
@@
[…]
@@ Appendix A Approximation of syntomic cohomology of classifying stacks
As in the proof of [ABM21, Theorem 1.2], we can find a $k$-linear representation $V$, a $d$-dimensional complete intersection $Z\subset\mathbb{P}(V)$ such that $Z$ is $G$-stable, $G$ acts freely on $Z$ and $Z/G\simeq[Z/G]$ is smooth and projective. For such a $Z$, [ABM21, Proposition 5.3 and 5.10] implies that the induced map $Z/G]\simeq Z/G\rightarrow[\mathbb{P}(V)/G]$ is a Hodge $d$-equivalence. Therefore, by Lemma A.2, the induced map on syntomic cohomology has cofibres in degrees $\geq d$. The claim then follows from the projective bundle formula in syntomic cohomology; we remark that this result for schemes is [BL22, Theorem 9.1.1] and the result follows by Kan extension for stacks which, in particular, shows that $H^{i}_{\mathrm{syn}}(\mathrm{B}G;\mathbb{F}_{p}(j))$ is a summand of $H^{i}_{\mathrm{syn}}([\mathbb{P}(V)/G];\mathbb{F}_{p}(j))$. Now, noting that $\mathrm{Spec}(k)\rightarrow\mathrm{B}G$ is an fpqc surjection, which pulls-back along $\mathbb{P}(V)/G]\rightarrow\mathrm{B}G$ to the fpqc surjection $\mathbb{P}(V)\rightarrow[\mathbb{P}(V)/G]$ the projective bundle formula follows from the one for schemes by the formula (37). ∎
@@ Appendix B Calculations in syntomic and motivic cohomology of certain classifying stacks
Throughout this section, we fix an perfect field $k$ of characteristic $p>0$. For an algebraic stack $\mathfrak{X}$ we define $\mathbb{F}_{p}(j)^{?}(\mathfrak{X})$ where $\mathrm{mot}$ or $\mathrm{syn}$ via right Kan extension from schemes. We write
@@
the bigraded cohomology ring of $k$ where $\mathrm{syn}$ or $\mathrm{mot}$. As already recorded in Example 4.11 we know how the cohomology of $B\mu_{p}^{\times s}$ takes the form:
@@
Here, |||-| refers to the topological/simplicial degree of a class/operation.
@@
We first give references for these equalities in the characteristic zero field $\mathbb{K}$. If $p$ is odd, then both equalities is [HW19, Lemma 13.11] (note that there should actually be a sign in the formulas because it is the case in topology and op. cit. obtained these formulas from the topological Steenrod algebra); note that in [HW19, Definition 13.10] the $\mathrm{Q}_{n}$’s are defined inductively by the right hand side of (1) and the cited reference proves the equality with our definition of $\mathrm{Q}_{n}$. If $p=2$, then we note that [Voe03b, Example 13.7] cautions that the commutator formula might not hold; it might also not be a derivation by [HW19, Remark 13.4.1]. However, the failure of both equalities are due to the presence of the element $\rho$ in motivic cohomology; such an element is zero if 1-1 is a sum of squares in the base field. We note that by examining [HW19, Corollary 13.14, Remark 13.14.1] both equalities hold in the case of $p=2$ modulo $\rho$. By the construction of our operations in characteristic $p>0$, both equalities now hold for all primes $p$ since $\rho$ is zero in characteristic $p>0$. The result for syntomic cohomology follows immediately. ∎